OpenStudy (anonymous):

A rotating beacon is located 7 miles out in the water. Let A be the point on the shore that is closest to the beacon. As the beacon rotates at 8 rev/min, the beam of light sweeps down the shore once each time it revolves. Assume that the shore is straight. How fast is the point where the beam hits the shore moving along the shore at an instant when the beam is lighting up a point 7 miles along the shore from the point A?

OpenStudy (anonymous):

I think it's going to be 7 * 8 = A

OpenStudy (anonymous):

I was thinking that it would be a triangle, and maybe applied derivatives of pythagarym theory?

OpenStudy (anonymous):

I think so too.......cuz I'm not good at problem solving but u could give me a medal for trying to answer ur question if u want

OpenStudy (turingtest):

@kludden you are very much on the right track

OpenStudy (anonymous):

Okay, I also looked at taking the derivative of tan(theta)= x/7 I'm not sure how to find out dx/dt though.

OpenStudy (turingtest):

dx/dt is what you are solving for what did you get after differentiating with respect to time the expression you wrote?

OpenStudy (anonymous):

for the tan equation or for the pythagaram theory?

OpenStudy (anonymous):

dx/dt should be the given rate that the light is travelling right?

OpenStudy (turingtest):

ye first of all, do you know d(theta)/dt ?

OpenStudy (anonymous):

no, I was thinking that would be what we were trying to find

OpenStudy (turingtest):

you can figure out d(theta)/dt from the fact that the thing is revolving at 8rev/min how many radians in one rev?

OpenStudy (anonymous):

If the beam moves in a circle 7 miles from shore, then youre looking at a circumfrence of 2pi*7. so 14pi. It moves 8 rev per minute, so 112pi ft/min.

OpenStudy (anonymous):

that could be incorrect

OpenStudy (anonymous):

2pi radians in a circle, so 10pi in 1 minute

OpenStudy (turingtest):

1revolutions is 2pi radians, so 8 rev/min=16pi rad/min=d(theta)/dt

OpenStudy (turingtest):

now differentiate\[\tan\theta=\frac x7\]with respect to time, what do you get?

OpenStudy (anonymous):

wow, yes that. My multiplication skills are superb lol. so 16pi rad in one minute. that gives you d(theta)/dt. You would get (1/7)*(1/(1+((x/7)^2))

OpenStudy (anonymous):

is equal to d(theta)/dt

OpenStudy (turingtest):

whoa, you differentiated wrong what is\[\frac d{dt}\tan\theta\]?

OpenStudy (anonymous):

sec^2 okay I took arctan to isolate theta but we have that so I was doing it wrong. you would get d(theta)/dt sec^2(theta)= (1/7) dx/dt

OpenStudy (turingtest):

yes, much better now just solve for dx/dt :)

OpenStudy (anonymous):

okay! Thank you so much!

OpenStudy (turingtest):

welcome!